OpenStudy (anonymous):

POLYGONS ABCD & EFGH ARE SIMILAR. WHAT IS THE PERIMETER OF EFGH WHEN AD = 12, BC = 8, FE = 15, FG = 10, and GH = 5?

OpenStudy (aum):



OpenStudy (aum):

Perimeter is the sum of the 4 sides. To find the perimeter of EFGH we need to find EH first. Since the two polygons are similar, the corresponding sides are proportional. Find the scale factor. BC abd FG are proportional. The scale factor is FG/BC = 10/8 = 5/4. The same scale factor applies to other sides as well. Therefore, EH/AD = 5/4 EH = 5/4 * AD = 5/4 * 12 = ? Then add up all the 4 sides of EFGH to find the perimeter.